# The Normal Play of the Domination Game

### v3

@@ Abstract
In 2010, Brešar, Klavžar and Rall introduced the optimization variant of the graph domination game and the game domination number. In 2024, Leo Versteegen obtained the celebrated proof of the Conjecture $\frac{3}{5}$ on this variant of the domination game, proposed by Kinnersley, West and Zamani in 2013. In this paper, we investigate for the first time the normal play of the domination game, which we call Normal Domination Game, that is an impartial game where the last to play wins. We first prove that this game is PSPACE-complete even in graphs with diameter two. We also use the Sprague-Grundy theory to prove that Alice (the first player) wins in the path $P_{n}$ if and only if $n$ is not a multiple of $4$, and wins in the cycle $C_{n}$ if and only if $n=4k+3$ for some integer $k$. Finally, we obtain a polynomial time algorithm to decide the winner for any disjoint union of paths and cycles in the Normal Domination Game and its natural partizan variant.
Keywords: Domination game, normal variant, PSPACE-hardness, Sprague-Grundy theory.
@@ 1 Introduction
The game variant of the independent set problem was introduced in 1908 by Henry Dudeney [5] as the Kayles game in which two players, Alice and Bob, alternately select vertices of a path graph $P_{n}$ in such a way that the set of selected vertices is an independent set. The Kayles game is in the normal play convention, that is, the last to play wins. In the misère play convention, the last to play loses and the game is called Misère Kayles. The Kayles game was solved in 1956 by Guy and Smith [7] through the Sprague-Grundy theory. The generalization of the Kayles game to any graph $G$ was denoted by Node Kayles game and was proved PSPACE-hard by Schaefer [11] in 1978. In 2002, Bodlander and Kratsch [2] proved that the Node Kayles game is polynomial time solvable in several graph classes, such as cographs, cocomparability graphs and circular arc graphs. In 2001, the optimization variant of the Node Kayles game was introduced [10], in which Alice wants to maximize the number of selected vertices and Bob wants to minimize this number. The game independence number $\alpha_{g}(G)$ is the maximum number of selected vertices in the game with best play for Alice and Bob. Curiously, this optimization variant of the Node Kayles game and the game independence number $\alpha_{g}(G)$ have been little studied in the literature.
The game variant of the dominating set problem was introduced much later, in 2010, by Brešar, Klavžar and Rall [4], called as the Domination Game. In this game, Alice and Bob alternately select vertices of a graph $G$ in such a way that every vertex selected in the game must dominate at least one vertex that was not dominated before. The game ends when the selected vertices form a dominating set. To the best of our knowledge, this game was investigated only in the optimization variant, in which Alice wants to minimize the number of selected vertices and Bob wants to maximize this number. The game domination number $\gamma_{g}(G)$ is the minimum number of selected vertices in the game with best play for Alice and Bob. This parameter was intensively investigated. In 2013, Kinnersley, West, and Zamani [9] proved the Theorem $2/3$ and proposed the Conjecture $3/5$ on the Domination Game. They state that $\gamma_{g}(G)$ is at most $\frac{2}{3}n$ and $\frac{3}{5}n$, respectively, if $G$ has no isolated vertices. Eleven years later, in 2024, Versteegen [13] obtained the celebrated proof of the Conjecture $3/5$, which is now called the Theorem $3/5$.
@@
In this paper, we start the investigation of the normal play of the domination game (the last to play wins), which we call Normal Domination Game. We first prove that the Normal Domination Game is PSPACE-complete. Our reduction is strongly based on the 1978 Schaefer’s PSPACE-hardness proof of the Node Kayles game [11].
[…]
@@ 2 Normal Domination game is PSPACE-hard
In this section, we prove the PSPACE-completeness of the Normal Domination game. We obtain a reduction from the Avoid-POSDNF-2 game, which was proved PSPACE-hard by Schaefer [11]. In this game, it is given a logical formula $\varphi$ in Disjunctive Normal Form [^1] Disjunction of conjunctions, such as $(x_{1}\land x_{4})\lor(x_{2}\land x_{4})\lor(x_{3}\land x_{4})\lor(x_{5}\land x_{6})$ (DNF) such that every conjunction has at most two variables and every literal is positive. All variables start false. Alice and Bob alternate turns assigning true to a variable valued false. The player that makes the formula true after their move loses. That is, the first player that assigns true to a variable that is in a conjunction alone or in a conjunction with other already chosen variable loses.
@@
In the Avoid-POSDNF-2 game, we may assume that every variable appears in at least one conjunction, no variable appears in all conjunctions and, for each variable, the number of conjunctions containing it is odd.
@@
If a variable $x_{i}$ does not appear in a conjunction, add the conjunction $(x_{i}\land x_{i}^{\prime})$, where $x_{i}^{\prime}$ is a new variable, and then $x_{i}$ or $x_{i}^{\prime}$ can be selected without interfering with the other conjunctions, but, if $x_{i}$ (resp. $x_{i}^{\prime}$) is selected when $x_{i}^{\prime}$ (resp. $x_{i}$) was already selected, the player loses. If some variable appears in all conjunctions, add a new variable $x$ and a new conjunction containing only $x$. With this, no variable appears in all conjunctions and playing in $x$ is losing. If some variable appears in an even number of conjunctions, we can make the following. Add a new variable $x^{\prime}$ and, for each variable that appears in an even number of conjunctions, add a conjunction with this variable and $x^{\prime}$. Moreover, add three (resp. two) conjunctions with $x^{\prime}$ alone if $x^{\prime}$ appears in an even (resp. odd) number of conjunctions. With this, each variable appears in an odd number of conjunctions. Moreover, no player wants to select $x^{\prime}$ (since otherwise they lose) and then the new conjunctions do not change the outcome of the game. ∎
@@
The Normal Domination Game is PSPACE-complete.
@@
Clearly, the problem is in PSPACE. Let $\varphi$ be an instance of the Avoid-POSDNF-2 game with variables $x_{1},\ldots,x_{n}$ and conjunctions $C_{1},\ldots,C_{m}$ satisfying Lemma 2.1. We may assume that $m\geq 3$, since otherwise the Avoid-POSDNF-2 game is easy. Let $G$ be the instance of the Normal Domination Game constructed in the following manner:
@@
  • Let $C=\{c_{1},\overline{c_{1}},c_{2},\overline{c_{2}},\ldots,c_{m},\overline{c_{m}}\}$ be a clique.
@@
Proof of Claim. Since the number of vertices in $V_{i}$ is odd and the number of unchosen vertices in $V_{i}$ is even, then there is an odd number of chosen vertices in $V_{i}$. This means that there is at least one vertex in $V_{i}$ already chosen at this point of the game, which implies that all vertices in $V_{i}$ are already dominated by the chosen vertices.
@@
Let $P_{1}$ be the player that selects a vertex in $V_{i}$ in the next turn and $P_{2}$ be the opponent. Since the game is open before $P_{1}$’s move, then no set $Q^{j}$ of size 1 contains a vertex of $V_{i}$ and every set $Q^{j}$ of size 2 containing a vertex of $V_{i}$ has a non-dominated vertex. Thus, after $P_{1}$’s move on $V_{i}$, the game remains open because it does not dominate any new vertex in $Q$. Therefore, $P_{1}$ cannot win with this move.
@@
After $P_{1}$’s move, there is an odd number of unchosen vertices in $V_{i}$. Let $v\in V_{i}$ be one such vertex, which is in some $Q^{j}$, and let $w$ be the other vertex in $Q^{j}$. Since the game is open and $v$ is dominated, $w$ is not dominated at this point. Thus, since no vertex of $C$ was chosen previously, $c_{j}$ is not dominated. Therefore, the opponent can play in $v$, since it dominates a new vertex ($c_{j}$). Also, the game is still open, since $v$ does not dominate any new vertex in $Q$. Finally, since before $P_{1}$’s turn, there was an even number of unchosen vertices in $V_{i}$ and both players selected vertices in $V_{i}$, then, after $P_{2}$’s turn, there is still an even number of unchosen vertices in $V_{i}$. $\diamond$
@@
Let $P_{1}$ be either Alice or Bob and let $P_{2}$ be the other player. Suppose that $P_{1}$ has a winning strategy in the Avoid-POSDNF-2 game. We show a winning strategy for $P_{1}$ in the Normal Domination Game. The strategy consists in choosing the corresponding vertices in $Q$ until the game closes. While the game is still open, no player wants to move in $C$; otherwise, the player loses from Claim 2.4. If $P_{2}$ chooses a vertex in some $V_{i}$ such that there is already a vertex in $V_{i}$ previously chosen, then, by Claim 2.6, $P_{1}$ can choose another vertex in $V_{i}$ so that the game remains open and the correspondence to the winning strategy in the Avoid-POSDNF-2 game is maintained. Since $P_{1}$ has a winning strategy in the Avoid-POSDNF-2 game, $P_{2}$ is the player that closes the game. Therefore, by Claim 2.5, $P_{1}$ wins the Normal Domination Game on $G$. ∎
@@ 3.1 Nimbers and the Sprague-Grundy Theory
Nim is one of the most important impartial games in which two players take turns removing objects from disjoint heaps until there are no objects remaining. On each turn, a player chooses a non empty heap and must remove a positive number of objects from this heap. An instance of Nim is given by a sequence $\Phi=(h_{1},h_{2},\ldots,h_{k})$, where $k$ is the number of heaps and $h_{i}\geq 1$ is the $i$-th heap’s size. Nim plays a fundamental role in the Sprague-Grundy game theory and was mathematically solved [3] by Bouton in 1901 for any instance $\Phi$. In other words, the decision problem of deciding which player has a winning strategy is polynomial time solvable in both Nim variants. This is related to nim-sum$(\Phi)=h_{1}\oplus\ldots\oplus h_{k}$, where $\oplus$ is the bitwise xor operation. Alice has a winning strategy in the normal play of Nim if and only if nim-sum$(\Phi)>0$. Moreover, Alice has a winning strategy in the misère play of Nim if and only if nim-sum$(\Phi)>0$ and there is at least one heap with more than one object, or nim-sum$(\Phi)=0$ and all non-empty heaps have exactly one object. The winning strategies in both variants is to finish every move with a nim-sum of 0, except in the misère variant when all non-empty heaps have exactly one object.
@@ 3.2 Paths
In this section, we prove the following theorem regarding the Normal Domination Game for paths $P_{n}$, where $P_{n}$ is the graph with vertices $n$ vertices $v_{1},\ldots,v_{n}$ and $n-1$ edges $v_{1}v_{2},v_{2}v_{3},\ldots,v_{n-1}v_{n}$. Let $n\ \%\ k$ be the remainder of the division of $n$ by $k$, also denoted by $n\mod k$.
@@
The proof is by induction on $n$. We first show the four base cases of $P_{n}^{\prime\prime}$ with $n=1,2,3,4$. Since $P_{1}^{\prime\prime}$ is a $P_{5}$ $v_{1}v_{2}v_{3}v_{4}v_{5}$ where $v_{1}$ and $v_{5}$ were already selected, the three possible moves ($v_{2},v_{3},v_{4}$) win immediately (nimber 00), and then the nimber of $P_{1}^{\prime\prime}$ is 1. Since $P_{2}^{\prime\prime}$ is a $P_{6}$ $v_{1}\ldots v_{6}$ where $v_{1}$ and $v_{6}$ were already selected, there are four possible moves ($v_{2}$ to $v_{5}$), which win immediately (nimber 00 in $v_{3}$ or $v_{4}$) or obtains a $P_{1}^{\prime\prime}$ (nimber 1 in $v_{2}$ or $v_{5}$), and then the nimber of $P_{2}^{\prime\prime}$ is $\operatorname{mex}\{0,1\}=2$. Since $P_{3}^{\prime\prime}$ is a $P_{7}$ $v_{1}\ldots v_{7}$ where $v_{1}$ and $v_{7}$ were already selected, there are five possible moves ($v_{2}$ to $v_{6}$), which win immediately (nimber 00 in $v_{4}$) or obtains a $P_{1}^{\prime\prime}$ (nimber 1 in $v_{3}$ or $v_{5}$) or obtains a $P_{2}^{\prime\prime}$ (nimber 2 in $v_{2}$ or $v_{6}$), and then the nimber of $P_{3}^{\prime\prime}$ is $\operatorname{mex}\{0,1,2\}=3$. Finally, since $P_{4}^{\prime\prime}$ is a $P_{8}$ $v_{1}\ldots v_{8}$ where $v_{1}$ and $v_{8}$ were already selected, there are six possible moves ($v_{2}$ to $v_{7}$), which obtains a $P_{1}^{\prime\prime}$ (nimber 1 in $v_{4}$ or $v_{5}$) or obtains a $P_{2}^{\prime\prime}$ (nimber 2 in $v_{3}$ or $v_{6}$) or obtains a $P_{3}^{\prime\prime}$ (nimber 3 in $v_{2}$ or $v_{7}$), and then the nimber of $P_{4}^{\prime\prime}$ is $\operatorname{mex}\{1,2,3\}=0$.
@@
Now we show the four base cases of $P_{n}^{\prime}$ with $n=1,2,3,4$. Since $P_{1}^{\prime}$ is a $P_{3}$ $v_{1}v_{2}v_{3}$ where $v_{1}$ was already selected in the game, the two possible moves ($v_{2}$ and $v_{3}$) wins immediately (nimber 00), and then the nimber of $P_{1}^{\prime}$ is 1. Since $P_{2}^{\prime}$ is a $P_{4}$ $v_{1}v_{2}v_{3}v_{4}$ where $v_{1}$ was already selected, there are three possible moves ($v_{2},v_{3},v_{4}$), which obtains a $P_{1}^{\prime}$ (with nimber 1 in $v_{2}$) or wins immediately (nimber 00 in $v_{3}$ or $v_{4}$), and then the nimber of $P_{2}^{\prime}$ is $\operatorname{mex}\{0,1\}=2$. Since $P_{3}^{\prime}$ is a $P_{5}$ $v_{1}v_{2}v_{3}v_{4}v_{5}$ where $v_{1}$ was already selected, there are four possible moves ($v_{2},v_{3},v_{4},v_{5}$), which obtains a $P_{2}^{\prime}$ (with nimber 2 in $v_{2}$) or obtains a $P_{1}^{\prime}$ (with nimber 1 in $v_{3}$) or wins immediately (nimber 00 in $v_{4}$) or obtains a $P_{1}^{\prime\prime}$ (with nimber 1 in $v_{5}$), and then the nimber of $P_{3}^{\prime}$ is $\operatorname{mex}\{2,1,0,1\}=3$. Finally, since $P_{4}^{\prime}$ is a $P_{6}$ $v_{1}v_{2}v_{3}v_{4}v_{5}v_{6}$ where $v_{1}$ was already selected, there are five possible moves ($v_{2},v_{3},v_{4},v_{5},v_{6}$), which obtains a $P_{3}^{\prime}$ (with nimber 3 in $v_{2}$) or obtains a $P_{2}^{\prime}$ (with nimber 2 in $v_{3}$) or obtains a $P_{1}^{\prime}$ (with nimber 1 in $v_{4}$) or obtains a $P_{1}^{\prime\prime}$ (with nimber 1 in $v_{5}$) or obtains a $P_{2}^{\prime\prime}$ (with nimber 2 in $v_{6}$), and then the nimber of $P_{4}^{\prime}$ is $\operatorname{mex}\{3,2,1,1,2\}=0$.

### v4

@@ Abstract
In 2010, Brešar, Klavžar and Rall introduced the optimization variant of the graph domination game and the game domination number, which was proved PSPACE-hard by Brešar et al. in 2016. In 2024, Leo Versteegen obtained the celebrated proof of the Conjecture $\frac{3}{5}$ on this variant of the domination game, proposed by Kinnersley, West and Zamani in 2013. In this paper, we investigate for the first time the normal play of the domination game, which we call Normal Domination Game, that is an impartial game where the last to play wins. We first prove that this game is PSPACE-complete even in graphs with diameter two. We also use the Sprague-Grundy theory to prove that Alice (the first player) wins in the path $P_{n}$ if and only if $n$ is not a multiple of $4$, and wins in the cycle $C_{n}$ if and only if $n=4k+3$ for some integer $k$. Moreover, we obtain a polynomial time algorithm to decide the winner for any disjoint union of paths and cycles in the Normal Domination Game and its natural partizan variant. Finally, we also prove that the Misère Domination Game (the last to play loses) is PSPACE-complete, as are the natural partizan variants of the normal game and the misère game.
Keywords: Domination game, normal game, misère game, PSPACE-hardness, Sprague-Grundy theory, partizan variant, Combinatorial Game Theory.
@@ 1 Introduction
The game variant of the independent set problem was introduced in 1908 by Henry Dudeney [6] as the Kayles game in which two players, Alice and Bob, alternately select vertices of a graph $G$ in such a way that the set of selected vertices is an independent set. The Kayles game is in the normal play convention, that is, the last to play wins. In the misère play convention, the last to play loses and the game is called Misère Kayles. The Kayles game was solved for some graph classes for the first time in 1956 by Guy and Smith [8] through the Sprague-Grundy theory. In 1978, Schaefer [12] proved that the Kayles game is PSPACE-hard. In 2002, Bodlander and Kratsch [2] proved that the Kayles game is polynomial time solvable in several graph classes, such as cographs, cocomparability graphs and circular arc graphs. In 2001, the optimization variant of the Kayles game was introduced [11], in which Alice wants to maximize the number of selected vertices and Bob wants to minimize this number. The game independence number $\alpha_{g}(G)$ is the maximum number of selected vertices in the game with best play for Alice and Bob. Curiously, this optimization variant of the Kayles game and the game independence number $\alpha_{g}(G)$ have been little studied in the literature.
The game variant of the dominating set problem was introduced much later, in 2010, by Brešar, Klavžar and Rall [4], called as the Domination Game. In this game, Alice and Bob alternately select playable vertices of a graph $G$, where a vertex is playable in the game if it dominates at least one vertex that was not dominated before. The game ends when the selected vertices form a dominating set. To the best of our knowledge, this game was investigated only in the optimization variant, in which Alice wants to minimize the number of selected vertices and Bob wants to maximize this number. The game domination number $\gamma_{g}(G)$ is the minimum number of selected vertices in the game with best play for Alice and Bob. This parameter was intensively investigated. In 2016, Brešar et al. proved that $\gamma_{g}(G)$ is PSPACE-hard [5]. In 2013, Kinnersley, West, and Zamani [10] proved the Theorem $2/3$ and proposed the Conjecture $3/5$ on the Domination Game. They state that $\gamma_{g}(G)$ is at most $\frac{2}{3}n$ and $\frac{3}{5}n$, respectively, if $G$ has no isolated vertices. Eleven years later, in 2024, Versteegen [14] obtained the celebrated proof of the Conjecture $3/5$, which is now called the Theorem $3/5$, in 33 pages.
@@
In this paper, we start the investigation of the normal play of the domination game (the last to play wins), which we call Normal Domination Game. We first prove that the Normal Domination Game is PSPACE-complete even in graph with diameter two. Our reduction is from the Avoid-POSDNF-2 game, proved PSPACE-hard by Schaefer [12]. We also use the Sprague-Grundy theory to prove that Alice (the first player) wins in the path $P_{n}$ if and only if $n\%4\neq 0$ ($n$ is not a multiple of $4$), and wins in the cycle $C_{n}$ if and only if $n\%4=3$, where $\%$ is the remainder operation (mod). We also determine in polynomial time the winner for any disjoint union of paths and cycles in the Normal Domination Game and its natural partizan variant, using the Combinatorial Game Theory. Finally, we also prove that the Misère Domination Game is PSPACE-complete, as are the natural partizan variants of the normal game and the misère game. We conclude with some open questions.
@@ 2 Normal Domination game is PSPACE-hard
In this section, we prove the PSPACE-completeness of the Normal Domination Game. We obtain a reduction from the Avoid-POSDNF-2 game, which was proved PSPACE-hard by Schaefer [12]. In this game, it is given a logical formula $\varphi$ in Disjunctive Normal Form [^1] Disjunction of conjunctions, such as $(x_{1}\land x_{4})\lor(x_{2}\land x_{4})\lor(x_{3}\land x_{4})\lor(x_{5}\land x_{6})$ (DNF) such that every conjunction has at most two variables and every literal is positive. All variables start false. Alice and Bob alternate turns assigning true to a variable valued false. The player that makes the formula true after their move loses. That is, the first player that assigns true to a variable that is in a conjunction alone or in a conjunction with another already chosen variable loses.
@@
Avoid-POSDNF-2 game is PSPACE-complete even if we assume that every variable appears in at least one conjunction, no variable appears in all conjunctions and, for each variable, the number of conjunctions containing it is odd.
@@
If a variable $x_{i}$ does not appear in a conjunction, add the conjunction $(x_{i}\land x_{i}^{\prime})$, where $x_{i}^{\prime}$ is a new variable, and then $x_{i}$ or $x_{i}^{\prime}$ can be selected without interfering with the other conjunctions, but, if $x_{i}$ (resp. $x_{i}^{\prime}$) is selected when $x_{i}^{\prime}$ (resp. $x_{i}$) was already selected, the player loses. If some variable appears in all conjunctions, add a new variable $x$ and a new conjunction containing only $x$. With this, no variable appears in all conjunctions and no player wants to select $x$, since otherwise they lose. If some variable appears in an even number of conjunctions, we can make the following. Add a new variable $x^{\prime}$ and, for each variable that appears in an even number of conjunctions, add a conjunction with this variable and $x^{\prime}$. Moreover, add three (resp. two) conjunctions with $x^{\prime}$ alone if $x^{\prime}$ appears in an even (resp. odd) number of conjunctions. With this, each variable appears in an odd number of conjunctions. Moreover, no player wants to select $x^{\prime}$ (since otherwise they lose) and then the new conjunctions do not change the outcome of the game. ∎
@@
The Normal Domination Game is PSPACE-complete even in graphs with diameter two.
@@
Clearly, the problem is in PSPACE. Let $\varphi$ be an instance of the Avoid-POSDNF-2 game with variables $x_{1},\ldots,x_{n}$ and conjunctions $C_{1},\ldots,C_{m}$ satisfying Lemma 2.1. We may assume that $m\geq 3$, since otherwise the Avoid-POSDNF-2 game is very easy. Let $G$ be the instance of the Normal Domination Game constructed in the following manner:
@@
  • Let $C=\{c_{1},\overline{c_{1}},c_{2},\overline{c_{2}},\ldots,c_{m},\overline{c_{m}}\}$ be a clique and let each $Q^{j}$ be an independent set.
@@
Proof of Claim. Since the number of vertices in $V_{i}$ is odd and the number of unchosen vertices in $V_{i}$ is even, there is an odd number of chosen vertices in $V_{i}$. This means that there is at least one vertex in $V_{i}$ already chosen at this point of the game, which implies that all vertices in $V_{i}$ are already dominated by the chosen vertices.
@@
Let $P_{1}$ be the player that selects a vertex in $V_{i}$ in the next turn and $P_{2}$ be the opponent of $P_{1}$. Since the game is open before $P_{1}$’s move, then no set $Q^{j}$ of size 1 contains a vertex of $V_{i}$ and every set $Q^{j}$ of size 2 containing a vertex of $V_{i}$ has a non-dominated vertex. Thus, after $P_{1}$’s move on $V_{i}$, the game remains open because it does not dominate any new vertex in $Q$. Therefore, $P_{1}$ cannot win with this move.
@@
After $P_{1}$’s move, there is an odd number of unchosen vertices in $V_{i}$. Let $v\in V_{i}$ be one such vertex, which is in some $Q^{j}$, and let $w$ be the other vertex in $Q^{j}$. Since the game is open and $v$ is dominated, $w$ is not dominated at this point. No vertex of $C$ was chosen previously, since, otherwise, $P_{1}$ could not have made its move on $V_{i}$. Thus, $c_{j}$ is not dominated. Therefore, the opponent can play in $v$, since it dominates a new vertex ($c_{j}$). Also, the game is still open, since $v$ does not dominate any new vertex in $Q$. Finally, since before $P_{1}$’s turn, there was an even number of unchosen vertices in $V_{i}$ and both players selected vertices in $V_{i}$, then, after $P_{2}$’s turn, there is still an even number of unchosen vertices in $V_{i}$. $\diamond$
@@
Let $P_{1}$ be either Alice or Bob and let $P_{2}$ be the other player. Suppose that $P_{1}$ has a winning strategy in the Avoid-POSDNF-2 game. We show a winning strategy for $P_{1}$ in the Normal Domination Game. The strategy consists in choosing the corresponding vertices in $Q$ until the game closes. While the game is still open, no player wants to move in $C$; otherwise, according to Claim 2.4, the player loses. If $P_{2}$ chooses a vertex in some $V_{i}$ such that there is already a vertex in $V_{i}$ previously chosen, then, by Claim 2.6, $P_{1}$ can choose another vertex in $V_{i}$ so that the game remains open and the correspondence to the winning strategy in the Avoid-POSDNF-2 game is maintained. Since $P_{1}$ has a winning strategy in the Avoid-POSDNF-2 game, $P_{2}$ is the player that closes the game. Therefore, by Claim 2.5, $P_{1}$ wins the Normal Domination Game on $G$. ∎
3 Misère Domination Game is PSPACE-hard
In this section, we prove the PSPACE-completeness of the Misère Domination Game, the variant of the domination game in which the last to play loses.
Theorem 3.1.
The Misère Domination Game is PSPACE-complete even in graphs with diameter four.
Proof.
Clearly, the problem is in PSPACE. We obtain a reduction from the normal variant. Let $H$ be an instance of the Normal Domination Game. Let $G$ be the graph obtained from $H$ by adding 6 new vertices $x_{1},x_{2},x_{3},y_{1},y_{2},y_{3}$ and the edges $x_{1}x_{2}$, $x_{1}x_{3}$, $y_{1}y_{2}$, $y_{1}y_{3}$, $x_{1}v$ and $y_{1}v$ for every vertex $v$ of $H$. See Figure 2.
$x_{1}$$y_{1}$$x_{2}$$y_{2}$$x_{3}$$y_{3}$HHGG
Figure 2: Proof of the PSPACE-hardness of the Misère Domination Game. Graph $G$ of the reduction from the Normal Domination Game on $H$.
Suppose that some player $X\in\{Alice,Bob\}$ has a winning strategy in the Normal Domination Game on $H$ and let $Y$ be the opponent. We present a winning strategy for $X$ in the Misère Domination Game on $G$. The player $X$ always prefer to play the winning strategy of the Normal Domination Game on the same vertices of $H$ (which are also vertices of $G$), unless $Y$ plays outside $H$. Since $X$ wins the normal variant in $H$, then $Y$ is the first to play outside $H$. If $y_{1}$ is the first vertex selected outside $H$, then $X$ selects $x_{2}$ and the only playable vertices are $x_{1}$ and $x_{3}$; thus $Y$ plays last and loses. If $y_{2}$ is the first vertex selected outside $H$, then $X$ selects $x_{1}$ and the only playable vertices are $y_{1}$ and $y_{3}$; thus $Y$ plays last and loses. Analogously if the first vertex selected outside $H$ is $y_{3}$, $x_{1}$, $x_{2}$ or $x_{3}$. ∎
@@ 4.1 Nimbers and the Sprague-Grundy Theory
[…]
@@ 4.2 Paths
In this section, we prove the following theorem regarding the Normal Domination Game for paths $P_{n}$, where $P_{n}$ is the graph with vertices $n$ vertices $v_{1},\ldots,v_{n}$ and n1n-1 edges $v_{1}v_{2},v_{2}v_{3},\ldots,v_{n-1}v_{n}$. Let $n\ \%\ k$ be the remainder of the division of $n$ by $k$, also denoted by $n\mod k$.
@@
The proof is by induction on $n$. We first show the four base cases of $P_{n}^{\prime\prime}$ with $n=1,2,3,4$. Since $P_{1}^{\prime\prime}$ is a $P_{5}$ $v_{1}v_{2}v_{3}v_{4}v_{5}$ where $v_{1}$ and $v_{5}$ were already selected, the three possible moves ($v_{2},v_{3},v_{4}$) win immediately (nimber 0), and then the nimber of $P_{1}^{\prime\prime}$ is 1. Since $P_{2}^{\prime\prime}$ is a $P_{6}$ $v_{1}\ldots v_{6}$ where $v_{1}$ and $v_{6}$ were already selected, there are four possible moves ($v_{2}$ to $v_{5}$), which win immediately (nimber 0 in $v_{3}$ or $v_{4}$) or obtains a $P_{1}^{\prime\prime}$ (nimber 1 in $v_{2}$ or $v_{5}$), and then the nimber of $P_{2}^{\prime\prime}$ is $\operatorname{mex}\{0,1\}=2$. Since $P_{3}^{\prime\prime}$ is a $P_{7}$ $v_{1}\ldots v_{7}$ where $v_{1}$ and $v_{7}$ were already selected, there are five possible moves ($v_{2}$ to $v_{6}$), which win immediately (nimber 0 in $v_{4}$) or obtains a $P_{1}^{\prime\prime}$ (nimber 1 in $v_{3}$ or $v_{5}$) or obtains a $P_{2}^{\prime\prime}$ (nimber 2 in $v_{2}$ or $v_{6}$), and then the nimber of $P_{3}^{\prime\prime}$ is $\operatorname{mex}\{0,1,2\}=3$. Finally, since $P_{4}^{\prime\prime}$ is a $P_{8}$ $v_{1}\ldots v_{8}$ where $v_{1}$ and $v_{8}$ were already selected, there are six possible moves ($v_{2}$ to $v_{7}$), which obtains a $P_{1}^{\prime\prime}$ (nimber 1 in $v_{4}$ or $v_{5}$) or obtains a $P_{2}^{\prime\prime}$ (nimber 2 in $v_{3}$ or $v_{6}$) or obtains a $P_{3}^{\prime\prime}$ (nimber 3 in $v_{2}$ or $v_{7}$), and then the nimber of $P_{4}^{\prime\prime}$ is $\operatorname{mex}\{1,2,3\}=0$.
@@
Now we show the four base cases of $P_{n}^{\prime}$ with $n=1,2,3,4$. Since $P_{1}^{\prime}$ is a $P_{3}$ $v_{1}v_{2}v_{3}$ where $v_{1}$ was already selected in the game, the two possible moves ($v_{2}$ and $v_{3}$) wins immediately (nimber 0), and then the nimber of $P_{1}^{\prime}$ is 1. Since $P_{2}^{\prime}$ is a $P_{4}$ $v_{1}v_{2}v_{3}v_{4}$ where $v_{1}$ was already selected, there are three possible moves ($v_{2},v_{3},v_{4}$), which obtains a $P_{1}^{\prime}$ (with nimber 1 in $v_{2}$) or wins immediately (nimber 0 in $v_{3}$ or $v_{4}$), and then the nimber of $P_{2}^{\prime}$ is $\operatorname{mex}\{0,1\}=2$. Since $P_{3}^{\prime}$ is a $P_{5}$ $v_{1}v_{2}v_{3}v_{4}v_{5}$ where $v_{1}$ was already selected, there are four possible moves ($v_{2},v_{3},v_{4},v_{5}$), which obtains a $P_{2}^{\prime}$ (with nimber 2 in $v_{2}$) or obtains a $P_{1}^{\prime}$ (with nimber 1 in $v_{3}$) or wins immediately (nimber 0 in $v_{4}$) or obtains a $P_{1}^{\prime\prime}$ (with nimber 1 in $v_{5}$), and then the nimber of $P_{3}^{\prime}$ is $\operatorname{mex}\{2,1,0,1\}=3$. Finally, since $P_{4}^{\prime}$ is a $P_{6}$ $v_{1}v_{2}v_{3}v_{4}v_{5}v_{6}$ where $v_{1}$ was already selected, there are five possible moves ($v_{2},v_{3},v_{4},v_{5},v_{6}$), which obtains a $P_{3}^{\prime}$ (with nimber 3 in $v_{2}$) or obtains a $P_{2}^{\prime}$ (with nimber 2 in $v_{3}$) or obtains a $P_{1}^{\prime}$ (with nimber 1 in $v_{4}$) or obtains a $P_{1}^{\prime\prime}$ (with nimber 1 in $v_{5}$) or obtains a $P_{2}^{\prime\prime}$ (with nimber 2 in $v_{6}$), and then the nimber of $P_{4}^{\prime}$ is $\operatorname{mex}\{3,2,1,1,2\}=0$.